After five games of bowling, Ralphs average score was 91. After six games his average score was 89. What was his score in the sixth game ?
Answer by TimothyLamb(4379)

x = sum of first five games' scores
y = score on sixth game
---
x/5 = 91
x = 91*5
---
(x + y)/6 = 89
x + y = 89*6
---
x + y = 89*6
91*5 + y = 89*6
y = 89*6 - 91*5
y = 79
---
answer:
y = score on sixth game = 79
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php